OpenStudy (anonymous):

lim sq rt. of x^2-9 lim as x approaches 16 a) sq. rt. 247 b) 123.5 c)+- sq rt. 247 d) does not exist

OpenStudy (anonymous):

\[\huge \lim_{x \rightarrow 16} {\sqrt{x^{2}-9}} = \sqrt{16^2-9} = ?\]

OpenStudy (nirmalnema):

option a is correct

OpenStudy (anonymous):

don;t just give answers, try leading them to it :D @nirmalnema

OpenStudy (nirmalnema):

you have given that so why will i give the same thing again